### v2

# From Simulations to Reality: Dark Energy Reconstruction with Simulated SNIa data from the Vera C. Rubin Observatory

Abstract
In this paper, we present an Artificial Neural Network (ANN) based reconstruction analysis of the Supernova Ia (SNIa) distance moduli ($\mu(z)$), and hence dark energy, using LSST simulated three-year SNIa data. Our ANN reconstruction architecture can model both the distance moduli and their corresponding error estimates. For this we employ astroANN and incorporate Monte Carlo dropout techniques to quantify uncertainties in our predictions. We tune our hyperparameters through advanced genetic algorithms, including elitism, utilizing the DEAP library. We compared the performance of the ANN based reconstruction with two theoretical descriptions of dark energy models, $\Lambda$CDM and Chevallier-Linder-Polarski (CPL). We perform a Bayesian analysis for these two theoretical models using the LSST simulations and also compare with observations from Pantheon and Pantheon+ SNIa real data. We show that our model-independent reconstruction using ANN is consistent with both of them. We assessed the performance using mean squared error (MSE) and showed that the ANN can produce distance estimates in better agreement with the LSST dataset than either $\Lambda$CDM or CPL, albeit very small. We included an additional residual analysis and a null test with $F$-scores to show that the reconstructed distances from the ANN model, are in excellent agreement with the $\Lambda$CDM or CPL model.
I Introduction
The discovery of the accelerated expansion of our Universe using Supernova Ia observations revolutionized our understanding of modern cosmology (riess_1998, ; Perlmutter:1998np, ; 2008ApJ…686..749K, ; Betoule2014, ; Planck2016, ; 2012ApJ…746…85S, ). The standard model of cosmology, $\Lambda$CDM, is composed of ordinary matter, a dark energy modeled as a cosmological constant $\Lambda$ that is responsible for the acceleration of the Universe, and Cold Dark Matter (CDM) Carroll:2000fy that shapes the cosmic structure through gravitational influence. $\Lambda$CDM assumes a homogeneous and isotropic Universe and has been in excellent agreement with most of the currently available data. However, it has its theoretical drawbacks such as fine-tuning and cosmic coincidence, and from an observational point of view, it also suffers from the Hubble tension (DiValentino:2021izs, ). These issues open the door to study models beyond the standard cosmological model.
There are non-parametric inference techniques that try to avoid assumptions of specific theoretical models, to find new dark energy properties from the data. They extract information directly from the data and infer unknown quantities based mainly on the data with as few assumptions as possible sahni2002cosmological ; wasserman2006all . Some interesting examples of these types of methods previously used to reconstruct cosmological functions are Principal Component Analysissharma2020reconstruction , smoothed step functions r8 , Gaussian processes williams2006gaussian ; Keeley:2020aym ; l2020defying ; r10 ; escamilla2023model and extrapolation methods montiel2014nonparametric . These approaches are also known as model-independent reconstructions, and they can be considered as new models based on the data, they can be used to analyze their similarity with different theoretical models and, to determine what model describes better the observational data. There are several works based on model-independent approaches to analyze dark energy features sahni2006reconstructing ; holsclaw2010nonparametric ; zhao2017dynamical ; r8 , cosmic expansion montiel2014nonparametric , deceleration parameter r10 , the growth rate of structure formation l2020defying ; gomez2023neuralrecs and luminosity distance wei2017improved ; lin2019non ; wang2020reconstructing .
In current times, the increase in computing power, bigger telescopes, enhanced CCDs, and a large amount of observational data have allowed the incorporation of machine learning methods as analysis tools in observational cosmology lin2017does ; peel2019distinguishing ; arjona2020can ; wang2020machine ; gomez2023neuralgenetic ; chacon2023analysis ; medel2023cosmological . Artificial Neural Networks (ANN) are one of the most revolutionary methods of Artificial Intelligence (AI), and they have been successfully employed in a wide range of applications in cosmology, for example, in image analysis dieleman2015rotation ; ntampaka2019deep , N-body simulations chacon2023analysis ; rodriguez2018fast ; he2019learning and statistical methods gomez2023neuralrecs ; gomez2023neuralgenetic ; alsing2019fast ; li2019model ; hortua2020constraining ; hortua2020parameter . From the data perspective, the forthcoming observational programs are anticipated to enhance the dark energy precision by almost ten folds FoM , some examples are Euclid EuclidTheoryWorkingGroup:2012gxx ; https://doi.org/10.48550/arxiv.0912.0914 , LSST LSSTDarkEnergyScience:2018jkl ; Zhan:2017uwu ; LSST:2008ijt , Roman Space Telescope Akeson:2019biv , the Thirty Meter Telescope (TMT) TMT , and the already operational JWST Gardner_2006 . Consequently, all the reasons above culminate in the increasing effort of leveraging AI and large datasets to investigate dark energy in a new way. This approach aims to overcome the longstanding issues of model degeneracy and the tension between different cosmological measurements without relying on predetermined theoretical models.
In this paper, we present a dark energy reconstruction analysis based on the state-of-the-art LSST’s three years of simulated PLAsTiCC SNIa data plasticc_announce and Artificial Neural Network modeling. We generated neural network models for the distance modulus and we compared this model with the Bayesian parameter estimation of dark energy models. In addition, we contrast the results with the SNIa data from the Pantheon and Pantheon+ surveys as real observational data. Our Bayesian analysis allows us to explore whether a model preference can be inferred from the LSST dataset, between the two most prevalent formulations of dark energy: $\Lambda$CDM and CPL (see §II.1). This paper is structured as follows, in sec. II we present the theory of the two different dark energy models we adopted in this analysis and a brief overview of the ANNs. In sec. III we outline the LSST survey program, which leads us to the next section IV, where we introduce the LSST simulated SNIa dataset; which is the primary dataset for our analysis. We also introduce the additional Pantheon SNIa datasets later in this section, used for comparisons only (and no ANN training is done with them). In the next sec. V, we discuss the methods used for constructing the ANN network and how the model is trained. In sec. VI we present the findings of our analysis and finally sec. VII presents the conclusion of this paper.
II Background
II.1 Dark Energy Models
We have considered the two most popular cosmological dark energy models, the $\Lambda$CDM and the Chevallier-Linder-Polarski (CPL) model (eos1, ; eos2, ), to test our LSST neural reconstruction in this analysis. The $\Lambda$CDM model often also called the concordance model, assumes that the Universe’s energy density is driven by a cosmological constant $\Lambda$ and it has a dark matter characterized as Cold Dark Matter (CDM). It is this cosmological constant that is believed to be due to dark energy. The dark energy component is parameterized by its Equation of State (EoS) parameter $w$ and is defined as:
$w=\frac{p}{\rho},$ (1)
and a cosmological constant $\Lambda$ corresponds to $w=-1$ and $\rho$ (energy density) constant. Over the last two decades, the $\Lambda$CDM model has consistently been shown to have good agreement with the observations. The Hubble parameter relates to the redshift ($z$) as,
$H^{2}(z)=H_{0}^{2}\left[\Omega_{m}(1+z)^{3}+(1-\Omega_{m})\right].$ (2)
where $\Omega_{m}$ is the matter density.
The second dark energy model we used in this analysis, is the CPL model. Parameterized as a function of scale factor $a$  [^1] Scale factor $a$ and redshift $z$ is linked as $a=(1+z)^{-1}$., it is a two-parameter dark energy model.
$w=w_{0}+w_{a}(1-a),$ (3)
it can be seen as a first-order approximation of the Taylor expansion of a more general dark energy EoS $w(a)$ (2017ApJ…850..183Z, ). This approach ensures that the model remains well-behaved across the entire range of the Universe’s history, from high to low redshifts. The dynamics of the Universe’s expansion are encapsulated by the scale factor $a$, with the background evolution described as,
$H^{2}(a)=H_{0}^{2}\left[\frac{\Omega_{m0}+a^{3}(1-\Omega_{m0})a^{-(3w_{0}+3w_{a}+3)}\exp\left(3w_{a}(a-1)\right)}{a^{3}}\right].$ (4)
The CPL is very well-behaved, all the way, from $a=0$ to $a=1$. Being a simple two-parameter model, the CPL is easy to analyze and also it adapts well to a wide variety of dark energy models.
II.2 Supernova Ia
Type Ia supernovae (SN Ia) are believed to be the result of the thermonuclear disruption of a carbon-oxygen white dwarfs which reaches the Chandrasekhar-mass limit of stability ($M_{Ch}\sim 1.4$M solar mass) by accreting matter from a companion 1960ApJ…132..565H . SNIa light curves show remarkable homogeneity after correcting for stretch and color parameters. Observed flux from SNIa can be used to compute the luminosity distance $D_{L}$. For an FLRW Universe, with a dark energy component EoS $w(z)$ the $D_{L}$ is given as,
$D_{L}=(1+z)\frac{c}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{E(z^{\prime})}$ (5)
where distance $D_{L}$ is in megaparsecs (Mpc) and $E(z^{\prime})$ is given as,
[…]
and the distance modulus $\mu$ can be related to the $D_{L}$ (for a flat Universe with a constant EoS dark energy $w=-1$) as
$\mu=5\log(D_{L}/10\text{pc}),$ (7)
therefore SNIa provides a direct and robust method to probe dark energy. Over the last two decades, numerous studies have utilized Type Ia Supernovae (SNIa) as the primary probe for investigating dark energy panstarrs ; Betoule2014 ; s1 ; Dhawan:2021ztf . Through successive SNIa programs, our precision in understanding dark energy has significantly improved, with efforts still ongoing Kessler2010 ; Linder2019 ; Mitra2021 .
II.3 Dark Energy Reconstruction and Artificial Neural Networks
Since the renewed interest in dark energy study, post the discovery of accelerating Universe via the SNIa observations (riess_1998, ; Perlmutter:1998np, ), the effort to reconstruct dark energy has been a very widely cultured topic sahni2006reconstructing ; re1 ; re2 ; re3 ; re4 ; re5 ; re6 ; re7 . Broadly, the reconstruction process can be classified either under parametric or non-parametric category. The parametric form of dark energy reconstruction has seen a lot of success, and it provides a lot of advantages, such as simplicity and interpretability, consistency in predictive power, especially in extrapolating beyond observed data, their simple forms allow easy efficiency of parametric models in fitting data. However parametric models also come with disadvantages. The biggest drawback of parametric models is the critical issues of model dependence and potential biases Liddle2007 . They also provide limited flexibility, in capturing unexpected features of the dark energy equation of state Copeland2006 , and finally, additional complexities can arise from parameter degeneracies, leading to over or underfitting Trotta2008 .
While parametric reconstruction methods have provided valuable insights by imposing specific models on the dark energy EoS $w(z)$, they inherently assume a certain level of theoretical bias toward the functional form of dark energy as discussed above. These limitations fanned the study of non-parametric reconstruction methods, which offer a more model-independent avenue for analyzing dark energy. Unlike their parametric counterparts, non-parametric methods do not rely on predefined equations or functions to describe the reconstructed parameter r8 ; r9 ; r10 ; r11 ; r12 . Instead, they leverage the data directly, allowing the underlying properties to manifest more freely in the analysis. In context to dark energy, this shift towards non-parametric reconstruction approaches can significantly open up new possibilities by reducing the theoretical prejudices and using the complexity and richness of the observational data r13 ; r14 ; gomez2023neuralgenetic . In this analysis, we use ANN-based non-parametric reconstruction technique gomez2023neuralrecs ; Wang:2020sxl ; Wang:2019vxv to obtain distance modulus ($\mu$) from SNIa observations. ANNs are powerful computational models capable of approximating any continuous nonlinear function, as demonstrated by hornik1990universal , making them exceptionally suited for modeling complex and large datasets. They are also better than the Gaussian Process (GP) based reconstructions in the sense that the ANN approach offers enhanced flexibility (eg. overfitting issues, kernel selection sensitivity, etc.) and accuracy in modeling natural processes compared to GP by having more optimizable hyperparameters. They are also less influenced by any prior assumptions which GP is more sensitive of dialektopoulos2022neural . For a comprehensive background on neural networks, seminal texts such as goodfellow2016deep ; bishop2006pattern ; nielsen2015neural provide in-depth analyses. The details of our ANN-based architecture are outlined in sec.§V.
III LSST: Overview
The Legacy Survey of Space and Time (LSST), conducted by the Vera C. Rubin Observatory collaboration, represents a paradigm shift in astrophysical surveys with its unparalleled scope and technical sophistication. Scheduled to begin in 2023, the LSST will deploy an advanced observational apparatus, featuring an $8.4$ m primary mirror with a $6.7$ m effective aperture and a state-of-the-art $3200$-megapixel camera, yielding a wide field of view of $9.6$ square degrees. The LSST is designed to survey approximately 18,000 square degrees of the southern sky over a decade, utilizing six optical passband filters to facilitate deep, wide, and fast observations. The survey aims to amass over 32 trillion observations of 20 billion galaxies and a similar number of stars, achieving a depth of $24^{th}$ magnitude in its six filter bands, spanning wavelengths from ultraviolet to near-infrared Cahn2009 ; ivezic . These efforts are projected to catalog millions of supernovae, among other transient phenomena, offering an unprecedented dataset for probing the dark Universe. The Dark Energy Science Collaboration (DESC)  [^2] https://lsstdesc.org/, comprising nearly 1,000 members, intends to harness this vast trove of data to extract high-precision measurements of fundamental cosmological parameters, leveraging prior data challenges to refine analysis pipelines in anticipation of the survey’s extensive data output (dc2, ; Sanchez, ).
IV Datasets
Our reconstruction analysis using the ANN models is based on the LSST 3-year simulated SNIa data. We have, also for comparison, analyzed the performance of the considered dark energy models using Pantheon and Pantheon+ SNIa datasets, all of which are described in this section below.
IV.1 Simulated LSST SNIa Data
We use three years of simulated Type-Ia Supernovae data from the LSST. This data is derived from mitra2023using and it consists of 5785 data points of SNIe with their simulated covariance matrix of statistical and systematic errors combined, within the redshifts $1.40.01<z<$ 1.4.
The SN data is generated using the LSST DESC time domain (TD) pipeline and SNANA code (snana, ), consisting of four main stages (illustrated in Fig. 2(a) of mitra2023using ). These include SN brightness standardization via a Light Curve (LC) fit stage, simulations for bias correction, and a BBC stage for Hubble diagram production before the last stage, cosmology fitting. We used a Hubble diagram and the associated covariance matrix (statistical + systematic) produced from the BBC stage to perform the cosmological fitting.
For mock generation, input cosmology: $\Omega_{m}=0.3150,\,\Omega_{\Lambda}=0.6850,\,w_{0}=-1,\,w_{a}=0$ was used. The curvature is computed internally as $\Omega_{k}=1-\Omega_{m}-\Omega_{\Lambda}$. It is based on cosmological parameters from Planck 2018 Planck2018 . In addition, the parameter $H_{0}$ is set to $70.0$ km/s/Mpc, this value is tied to SALT2 training salt and we use the SALT2 lightcurve model. It generates observer frame magnitudes. The noise in the simulation is computed as follows :
$\sigma_{\text{SIM}}^{2}=\left[F+(A\cdot b)+(F\cdot\sigma_{\text{ZPT}})^{2}\right.\left.+\sigma_{0}\cdot 10^{0.4\cdot ZPT_{\text{pe}}}+\sigma_{\text{host}}^{2}\right]S^{2}_{\text{SNR}},$ (8)
where $F$ is the simulated flux in photoelectron (p.e.), $A$ is the noise equivalent area given by, $A=\left[2\pi\int\text{PSF}^{2}(r,\theta)r\mathop{}\!\mathrm{d}r\right]^{-1}$ where PSF stands for the Point Spread Function, $b$ is the background per unit area (includes sky + CCD readouts + dark current), $S_{\text{SNR}}$ is an empirically determined scale that depends on the signal-to-noise ratio, the three $\sigma$ terms correspond to zero point uncertainty, flux calibration uncertainty and underlying host galaxy uncertainty. These terms can be determined empirically from fits that match simulated uncertainties to those from the survey designs.
We use a redshift-binned Hubble diagram (HD) alongside its corresponding covariance matrix. The datasets are composed of a mixture of spectroscopically ($z_{\rm spec}$) and photometrically ($z_{\rm phot}$) identified SNIa candidates at low redshift and high redshift respectively.
The photometric redshift determination and its uncertainty were based on Graham2018_photoz while using host galaxy photo-$z$ as priors (Kessler2010 ). mitra2023using re-simulated the PLAsTiCC  data based on the DDF strategy of the LSST, augmented with low-redshift spectroscopic data from the DC2 analysis. The study by Mitra:2022ykq , which simulated exclusively SNIa without considering contamination, generated a statistical plus systematic covariance matrix to account for seven systematics. The HD is comprised of 5809 SNIa candidates. Importantly, the work by Mitra:2022ykq expanded on the initial analysis plan for the LSST outlined by LSSTDarkEnergyScience:2018jkl , which focused on a 1 and 10-year timeline for studying Supernova Type Ia (SNIa) cosmology. Mitra:2022ykq did this by conducting a comprehensive analysis that takes into account both the spectroscopic and photometric redshifts of the supernova’s host galaxies.
IV.2 Pantheon and Pantheon+ compilations
The Pantheon datasets are all publicly available in the form of unbinned Hubble diagram data. Pantheon has 1048 data points scolnic2018complete within redshifts between $z=0.01$ and $z=2.3$, and Pantheon+ has 1550 Type Ia between $z=0.001$ and $z=2.26$ brout2022pantheon+ . They are the latest compilations of the SNIa observations, expanding from the previously concluded confirmed Pan-STARRS1 survey 2002SPIE.4836..154K . Unlike the LSST simulated data, all of Pantheon’s data are spectroscopic in nature.
V Methodology
This section describes how we use the ANN to generate neural network reconstructions for the distance modulus. We employ the methodology presented in gomez2023neuralrecs using the hyperparameter tuning suggested in gomez2023neuralgenetic . In our case, we use the TensorFlow library to implement our neural network model. Below we outline the steps:
  • We prepare the dataset for the training of our neural network. We consider the redshift ($z$) as input, and the distance modulus ($\mu$) with the error ($\sigma(\mu)$) conformed by the sum of their statistical and systematic errors.
  • We explore the dataset and train different architectures of neural networks with Monte Carlo Dropout gal2015dropout (as in gomez2023neuralrecs ) to gain empirical knowledge about the range of hyperparameters. The Monte Carlo Dropout allows to obtain the uncertainties of the neural network predictions and to generate a more robust model based on the data. We use the Monte Carlo dropout implementation from AstroANN astronnleung2019deep ; astronn .
  • To find a proper combination of hyperparameters, we set the search space for the use of nnogada (gomez2023neuralgenetic, ) and their implemented simple genetic algorithms with elitism from the DEAP library DEAP_JMLR2012 ; de2014deap (for details on genetic algorithms we recommend Ref. medel2023cosmological ). In our case, we consolidated the search space with the number of layers $\in[3,4]$, the number of neurons for hidden layers $\in[100,200]$, and batch size $\in[8,16,32,64]$.
  • The best combination found was 4 layers with 200 neurons in each one and a batch size of 32 with a loss function value of 0.02528 (see Fig. 1).
  • We use the best combination to build our ANN network and apply the training data to it.
  • Once the neural network is trained, we perform several predictions of the distance modulus to generate a non-parametric reconstruction based on the LSST simulated SNIa data.
  • We test with the two dark energy models discussed previously. We compare the theoretical predictions and the reconstruction with the data, and we measure with the MSE metric to test which cases are in more agreement with the dataset. The values of the dark energy model parameters are previously fitted with Bayesian inference using the same data used for the neural reconstruction.
Figure 1: ANN architecture founded by the hyperparameter tuning with genetic algorithms using nnogada.
VI Results & Discussions
VI.1 Bayesian Analysis
To have an insight into the cosmological models considered in this work constrained by the LSST data, we perform an initial Bayesian inference using nested sampling with the dynesty library speagle2020dynesty and the SimpleMC code aubourg2015 ; simplemc . We also do it with Pantheon and Pantheon+, to compare the results obtained with the LSST simulations. This provides us with a baseline reference, before the subsequent ANN analysis.
Table 1 includes the ranges of the priors used for the free parameters of both $\Lambda$CDM and CPL. In addition, it shows the results of posterior distribution sampling and the Bayes factor to compare models. According to Fig. 2 and Table 1, we can notice that for $\Lambda$CDM the $\Omega_{m}$ is considerably better constrained with the LSST data than the other two SNIa compilations; while the $h$ parameter is practically not affected between the three different datasets. On the other hand, for the CPL model, the parameter estimation with LSST data is consistent with Pantheon and Pantheon+, with a better constraint for the $w_{0}$ parameter.
In all the cases the $-2\log\mathcal{L}_{max}$ (Table 1) obtains a lower value for the LSST dataset, which implies that this data obtains a better fit for the cosmological models. In addition, something interesting is that using the LSST data, the Bayesian model comparison suggests that $\Lambda$CDM increases its advantage over CPL because according to Jeffrey’s scale, both Pantheon and Pantheon+ $\Lambda$CDM only have weak evidence, and for LSST data this scale turns to moderate. This disadvantage can also be analyzed in Fig. 3, where the EoS for CPL using LSST data puts the $w=-1$ line for $\Lambda$CDM almost out of $1-\sigma$, while for Pantheon and Pantheon+, the most probable value for $w$ is statistically closer to the $\Lambda$CDM prediction. However, considering only the values of $-2\log\mathcal{L}_{max}$, for the three datasets, CPL has always a better fit to the data than $\Lambda$CDM; therefore, the Bayesian evidence could be very sensitive to the extra-parameters of CPL and, in future work, it could be worth to incorporate other model comparison techniques because Bayesian evidence has received some criticisms in cosmological data analysis efstathiou2008limitations ; nesseris2004comparison ; koo2022bayesian . However, our results clearly show that the LSST data can give more advantages while investigating for model selection tests in the future (more so with LSST $10$ years SNIa data) though in the current analysis our results from Table 1, are insufficient and very weak to choose in favor of any model.
Priors $\Lambda$CDM Pantheon CPL Pantheon $\Lambda$CDM Pantheon+ CPL Pantheon + $\Lambda$CDM LSST CPL LSST
$\Omega_{m}$ [0.05, 0.5] $0.3017\pm 0.0220$ $0.3203\pm 0.0821$ $0.3317\pm 0.0184$ $0.2846\pm 0.0930$ $0.3163\pm 0.0049$ $0.3297\pm 0.0700$
$h$ [0.4, 1.0] $0.7057\pm 0.1755$ $0.6430\pm 0.1427$ $0.6837\pm 0.1690$ $0.7001\pm 0.1667$ $0.6967\pm 0.1732$ $0.7125\pm 0.1585$
$w_{0}$ [-2.0, 0.0] - $-1.0960\pm 0.2087$ $-0.9218\pm 0.1372$ $-1.0168\pm 0.0679$
$w_{a}$ [-2.0, 2.0] - $-0.0996\pm 1.0031$ $-0.0666\pm 0.8226$ $-0.4305\pm 0.8789$
$-2\log\mathcal{L}_{max}$ 1024.9833 1024.9165 1403.1125 1402.6494 5502.7526 5501.2499
$\log Z$515.9625 -517.4652 -705.2655 -706.8580 -2756.3249 -2759.0397
$\log B_{\Lambda CDM,CPL}$ 1.5027 - 1.5925 - 2.7148
Evidence - - Weak - Weak - Moderate
Table 1: Bayesian analysis to $\Lambda$CDM and CPL models using Pantheon, Pantheon+ and LSST data. We can notice the priors of their parameters, their parameter estimation, $-2\log\mathcal{L}_{max}$, the log-Bayesian evidence $\log Z$, the log-Bayes factor $\log B$ of CPL and $\Lambda$CDM for each dataset. Log-Bayes factor is defined as $\ln B_{ab}\equiv\ln Z_{a}-\ln Z_{b}$, and according to Jeffrey’s scale wasserman2000bayesian we can consider weak evidence if $10\leq|\Delta\log Z|<$ 1, moderate evidence if $31\leq|\Delta\log Z|<$ 3, strong evidence if $53\leq|\Delta\ln Z|<$ 5, and decisive evidence if $|\Delta\log\mathcal{Z}|\geq 5$, in favourof the preferred model; if $\Delta\log Z$ is negative it is in favourof the second term, and if it is positive the first term is favoured hee2016bayesian . In our results, $\Lambda$CDM is always the preferred model.
(a) $\Lambda$CDM
(b) CPL
Figure 2: One and two-dimensional posterior distributions from Bayesian sampling for the free parameters of $\Lambda$CDM (left) and CPL (right) using SNIa data from Pantheon, Pantheon+, and LSST simulations.
(a) Pantheon
(b) Pantheon+
(c) LSST
Figure 3: EoS for CPL with the three different datasets. We obtained these plots from the posterior distribution sampling using fgivenx handleyfgivenx . It is interesting to note that for the LSST case, the $w=-1$ line (red dashed) marginally crosses the $1-\sigma$ range.
VI.2 Neural Reconstruction
Finally, after careful training of a neural network model with the LSST simulated data, we generated the model-independent reconstruction for $\mu(z)$ as can be seen in Fig. 4. In Fig.4 we plot the distance modulus ($\mu$) from the datasets of LSST sim and Pantheon, Pantheon+, and then we overlay with the reconstructed distance modulus (black dashed lines) along with their associated uncertainties from the ANN predictions plus the modeled errors (yellow regions). We also plot the theoretically computed $\mu(z)$ from the two cosmological dark energy models of $\Lambda$CDM and CPL. From the left panel of Fig. 4 it can be seen that the reconstructed data points from the ANN model and the LSST observations (green points) are very close. Next, we computed the MSE for each of these three cases with the observed LSST dataset, shown in Table 2. From Table 2 we see that MSE for ANN is slightly better than $\Lambda$CDM or CPL. This implies that the artificial neural network (ANN) based non-parametric reconstruction model can indeed more closely replicate the data than the theoretical models, albeit by a small margin. The second row in Table 2 gives additional insight by comparing the MSE between the ANN reconstructed $\mu(z)$ and the corresponding $\Lambda$CDM or CPL based distances. It is worth noting that, due to the hyperparameter tuning of the ANN and the Monte Carlo dropout, we have a good model based on the data, which looks visually smooth and does not have any overfitting or underfitting in machine learning terms.
In Fig. 5 we present residual plots for the distance modulus $\mu(z)$, computed as $\delta\mu\mid_{i}=\mu_{i}-\mu_{model}(z_{i})$ (where $\mu_{i}$ corresponds to the $i^{th}$ observation from the LSST sim data) for two cases, a) LSST sim data - $\Lambda$CDM theory and b) LSST sim data - ANN model reconstructed $\mu(z)$ (we compare the $\mu(z)$ residuals between the ANN model and the $\Lambda$CDM model only, as $\Lambda$CDM is seen to have bigger differences with the ANN reconstruction as seen in Table 2). It can be seen that there is an infinitesimal difference between the residuals from $\Lambda$CDM and ANN model. However, we can notice the sensitivity of the ANN model from outliers, which can make the error margins get bigger. We performed an additional $F$-test to see if there is any statistically significant difference between the distance modulus residuals from $\Lambda$CDM and ANN model  [^3] The $F$-test is used to compare the variances of two independent samples to test the hypothesis that they are equal. It is defined as the ratio of the variances, $F=\frac{\sigma_{1}^{2}}{\sigma_{2}^{2}}$, where $s_{1}^{2}$ and $s_{2}^{2}$ are the sample variances. This statistic follows the $F$-distribution under the null hypothesis that the variances are equal snedecor1989statistical . From the $F$-test score, the corresponding $p$ value computed is $p\simeq 0$, thereby signifying that the ANN model does not provide any statistically significant improvement over the $\Lambda$CDM estimates.
Refer to caption Refer to caption
Figure 4: Neural network reconstruction for distance modulus, $\mu(z)$ (black) with the standard deviation of the predictions plus the modeled error (yellow region) using LSST data (green dots). Left: With $\Lambda$CDM and CPL using the values of the parameter estimation using LSST data. Right: In comparison with data points from other SNIa surveys, Pantheon and Pantheon+ using $\Lambda$CDM model only.
$\Lambda$CDM CPL ANN
LSST 0.05164 0.04708 0.03797
Neural predictions 0.02051 0.01881
Table 2: Mean squared error between $\mu(z)$ predictions for (first row) three different models ($\Lambda$CDM, CPL and ANN) with the observed $\mu(z)$ from the LSST SNIa sim dataset and (second row) between the ANN reconstructed distance estimates with the corresponding theoretical estimates from $\Lambda$CDM and CPL.
Figure 5: Residual plot ($\delta\mu$) for two cosmological models as shown in figure legend. The residuals are computed between the LSST SNIa sim data (Fig. 4) and model-derived distance estimates. The right panel shows a Gaussian distribution of the corresponding residuals.
VII Conclusions
In this paper, we have provided a proof of concept, to forecast a deep learning-based non-parametric dark energy reconstruction model. Which uses the state of the art deep learning tools to find solutions to the open question of finding the most accurate dark energy model. While trying to maximize the benefits one can draw from the huge SNIa datasets that the LSST, is going to provide in a couple of years. This paper gives a flavour of how a deep learning based dark energy reconstruction model, for SNIa observations, can be ideally implemented with the actual observation of SNIa in the years to come, not only for the LSST but also from other similar programs (with suitable modifications). In this paper, we show unequivocally that our reconstructed ANN-based model does perform (slightly) better than the traditional $\Lambda$CDM and CPL models when tracing the observational data. But it also shows possible areas for future studies. For example, the increased standard deviation observed in the ANN model compared to the theoretical model (Fig. 5) underscores the need for further investigation. This work not only demonstrates the viability of deep learning for cosmological reconstruction analysis but also opens avenues for future research to refine these models for higher precision in understanding dark energy better.
In addition, we notice that the $\Lambda$CDM model is favoured in comparison to CPL using the LSST data, suggesting that there might be room for exploring alternate parameterizations of the dark energy equation of state. In contrast, the MSE of the CPL considering the neural reconstruction is better than $\Lambda$CDM, and considering that despite the lack of interpretability, our neural reconstructions have been rigorously trained to have a generalizable model based only on the data (without underfitting and overfitting), without any cosmological or statistical assumptions, therefore in this CPL shows an advantage over $\Lambda$CDM. These results suggest that it is necessary to incorporate other model comparison techniques to have greater confidence in the preference of the data over the theoretical models, and the nonparametric models may be an opportunity to expand the current repertoire.
We can highlight that the forthcoming LSST data may improve the constraints on parameters of the dark energy models, as we have shown in our Bayesian data analysis. On the other hand, using the parameter estimation for $\Lambda$CDM and CPL with LSST data and comparing it with the neural network reconstruction based on LSST simulations, we have shown that there is another model, smooth and without overfitting, that agrees better with these SNIa data. This indicates that further exploration of new theoretical models together with new statistical and computational models based on the forthcoming cosmological datasets will help to have a broader picture of the directions to follow to understand the behaviour of dark energy, presumably with more machine learning involvement.
Acknowledgements.
AM acknowledges useful discussions with Surhud More and Purba Mukherjee. AM also acknowledges Richard Kessler and the LSST DESC for generating the LSST SNIa simulated dataset. IGV thanks the CONACYT postdoctoral grant.